OpenStudy (codysek98):

Find the absolute maximum and minimum values of the function on the given interval. Then graph the function. Identify the points on the graph where the absolute extrema occur. I'll start typing the equation below in a reply.

OpenStudy (codysek98):

\[f(\theta) = \sin(\theta), -(\frac{ 5\pi }{ 3 }) \le \theta \le-(\frac{ \pi }{ 2 })\]

OpenStudy (solomonzelman):

\(f'(\theta)={\tiny~}?\)

OpenStudy (codysek98):

that's the derivative of f(theta), right?

OpenStudy (solomonzelman):

Yes.

OpenStudy (codysek98):

f'(theta) = cos(theta) is that correct?

OpenStudy (solomonzelman):

Yes, you are correct!

OpenStudy (codysek98):

do i plug in -5pi/3 for the min and plug in -pi/2 for the max?

OpenStudy (solomonzelman):

So, the possible critical numbers, if the interval for \(\theta\) is given are: 1. Interval boundaries (only closed boundaries, not open boundaries). 2. All values (within the given interval) where \(f'(\theta)\) is undefined (provided \(f(\theta)\) is defined at that value.

OpenStudy (solomonzelman):

So, yes, you plug in the interval boundaries.

OpenStudy (codysek98):

cos(5pi/3) = .5 is that right?

OpenStudy (solomonzelman):

also, 3. Any points where \(f'(\theta)=0\).

OpenStudy (solomonzelman):

yes, you are right.

OpenStudy (codysek98):

and then cos(-pi/2) = 0

OpenStudy (solomonzelman):

Yes

OpenStudy (solomonzelman):

And sin(-pi/2)=1, so you definitely have a local minimum at theta=-pi/2

OpenStudy (solomonzelman):

BUT, you plug in the boundaries into the function, not into the derivative.

OpenStudy (codysek98):

oh okay. and how did you know for sure that there was a min at -pi/2 just because it equaled 1?

OpenStudy (solomonzelman):

Oh, I meant sin(-pi/2)=-1

OpenStudy (solomonzelman):

not positive 1, sorry. And why did I know that? Because sine of any angle is always AT MOST +1, and AT LEAST -1.

OpenStudy (solomonzelman):

In other words, \(-1\le\sin[f(\theta)]\le1\).

OpenStudy (codysek98):

oh okay. so the absolute minimum is -1 at -pi/2

OpenStudy (solomonzelman):

Yes, \(f(\theta)=-1\), at \(\theta=-\pi/2\).

OpenStudy (codysek98):

i mean absolute minimum is -1 at theta = -pi/2

OpenStudy (solomonzelman):

yes, yes.

OpenStudy (codysek98):

okay. so for the maximum, i now know the max would be positive one at theta = -5pi/3

OpenStudy (solomonzelman):

Actually, you are not correct.

OpenStudy (codysek98):

oh wait is it .5?

OpenStudy (solomonzelman):

\(\sin\theta\) has a period every \(2\pi\). So, if it has a minimum at \(\theta=-\pi/2\), then it has other closest minimums at \(\theta=3\pi/2\) and \(\theta=-5\pi/2\).

OpenStudy (solomonzelman):

maximums of any sine function are right in between the minimums.

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/kxfbopjxxp

OpenStudy (codysek98):

so would the maximum be 1 at theta = -3pi/2 ?

OpenStudy (solomonzelman):

Yes.

OpenStudy (codysek98):

thank you